OpenStudy (anonymous):

identify the system of equations. 3x + 2y = 7 -4x- 3y = 10

OpenStudy (stormfire1):

3x+2y=7 -4x-3y=10 Add the equations together: 3x-4x+2y-3y = 7 + 10 -x-y = 7 + 10 So I believe this is an independent with many solutions.